# Fast Frank–Wolfe Algorithms with Adaptive Bregman Step-Size for Weakly Convex Functions

### v3

@@ A.2 Lemmas of Smooth Adaptable Property
Lemma A.4 (Extended Descent Lemma [Bolte2018-zt, Lemma 2.1]).
@@
From this, it can be seen that the $L$-smad property for $(f,\phi)$ provides upper and lower approximations for $f$ majorized by $\phi$ with $L>0$. In addition, if $\phi(x)=\frac{1}{2}\|x\|^{2}$ on $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$, Lemma A.4 corresponds to the classical descent lemma. While the $L$-smad property might seem unfamiliar at first, it is a natural generalization of $L$-smoothness, and examples of functions $f$ and $\phi$ satisfying the $L$-smad property are given, e.g., in [Bauschke2017-hg, Lemmas 7 and 8][Bolte2018-zt, Lemma 5.1], [Dragomir2021-rv, Propositions 2.1 and 2.3], [Mukkamala2019-mk, Proposition 2.1], [Takahashi2024-wz, Proposition 24], [Takahashi2023-uh, Theorem 1], and [Takahashi2024-cu, Theorem 15].
@@
Lemma A.5 (Primal progress from the smooth adaptable property).
@@
Using Lemma A.4 and substituting $x_{+}$ for $x$ and $x$ for $y$, we have
@@
It holds that $D_{\phi}(x_{+},x)\leq\gamma^{1+\nu}D_{\phi}(v,x)$ for some $\nu>0$ by Lemma A.3. Therefore, this provides the desired inequality. ∎
@@
Lemma A.6 (Progress lemma from the Bregman short step-size).
@@
Substituting $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$ with $x_{t},v_{t}\in P$ for $x_{+}$ in Lemma 2.4, we have
@@
Appendix B Properties from Sharpness
B.1 Results from Weak Convexity
Obviously, $f$ is $\rho$-weakly convex and differentiable if and only if, for all $x,y\in\operatorname{dom}f$,
@@ B.1 Results from Weak Convexity
Any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex [Vial1983-oo, Proposition 4.11] and [Wu2007-dt]. Since the proof of this fact is omitted therein, we provide the argument for the sake of completeness below.
Proposition B.1.
@@
Let $g=\lambda_{\min}(\nabla^{2}f(\cdot))$, where $\lambda_{\min}(\cdot)$ is the smallest eigenvalue. Obviously, $g$ is continuous because $\lambda_{\min}(\cdot)$ and $\nabla^{2}f(\cdot)$ are continuous. Therefore, there exists the minimum value of $g$ on $P$ due to the continuity of $g$ and the compactness of $P$. For $\rho=|\min_{x\in P}g(x)|$, we have $\nabla^{2}f(x)+\rho I\succeq O$ for any $x\in P$, which implies $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex on $P$, i.e., $f$ is $\rho$-weakly convex on $P$. ∎
Minimizing a nonconvex $\mathcal{C}^{2}$-function over a compact set $P$ is equivalent to minimizing a weakly convex function over $P$. However, it is important to note that when $\rho$ becomes large, the assumptions of algorithms may not be satisfied (see also Theorems D.3 and 5.3).
@@
Lemma B.2 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex functions).
Let $x^{*}\in P$ an optimal solution of (1.1) and let $f^{*}=f(x^{*})$. Suppose that Assumptions 3.1iiiv, and that $f$ is $\rho$-weakly convex. Then for all $x\in P$, it holds:
$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (B.1)
@@
B.2 Primal Gap Bound
We show in the following lemma that the HEB condition immediately provides a bound on the primal optimality gap, which will be useful for establishing convergence rates of the proposed algorithms as well as the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks], provided $f$ is convex.
Lemma B.3 (Primal gap bound from Hölder error bound).
@@ B.2 Primal Gap Bound
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}},$ (B.2)
@@
Remark B.4.
With the remark from above, we can immediately relate the HEB condition to the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks]; the Łojasiewicz inequality is used to establish convergence analysis [Bolte2007-xr]. To this end, using (B.2), we estimate $\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\leq\frac{\|\nabla f(x)\|\|x-x^{*}\|}{\|x-x^{*}\|}=\|\nabla f(x)\|$ to then obtain the weaker condition:
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\|\nabla f(x)\|^{\frac{q}{q-1}}=c^{\frac{1}{\theta}}\|\nabla f(x)\|^{\frac{1}{\theta}},$ (B.3)
where $c=\left(\frac{q}{\mu}\right)^{\frac{1}{q}}$ and $\theta=\frac{q-1}{q}$. Inequality (B.3) is called the $c$-Łojasiewicz inequality with $\theta\in[0,1)$. If $\mathcal{X}^{*}\subseteq\operatorname{ri}P$, then the two conditions are equivalent. However, if the optimal solution(s) are on the boundary of $P$ as is not infrequently the case, then the two conditions are not equivalent as $\|\nabla f(x)\|$ might not vanish for $x\in\mathcal{X}^{*}$, whereas $\langle\nabla f(x),x-x^{*}\rangle$ does, i.e., the HEB condition is tighter than the one induced by the Łojasiewicz inequality.
@@
Lemma B.5 (Primal gap bound from the quadratic growth).
Let $f$ be a $\rho$-weakly convex function that satisfies the local $\mu$-quadratic growth condition such that $\rho\leq\mu$. Let $x^{*}$ be the unique minimizer of $f$ over $P$ and let $\zeta>0$. Then, the following holds: for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$ (B.4)
@@
$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\langle\nabla f(x),x-x^{*}\rangle=\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\|x-x^{*}\|.$ (B.5)
@@
Remark B.6.
In the same vein as the discussion in Remark B.4, we can immediately relate the local $\mu$-quadratic growth condition with $\zeta>0$ to the local Polyak–Łojasiewicz (PL) inequality. Using (B.4), for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\|\nabla f(x)\|^{2},$ (B.6)
which is equivalent to the PL inequality [Polyak1963-dr], also called the gradient dominance property [Combettes2020-jj], with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$. The PL inequality is also equivalent to the Łojasiewicz inequality with $\theta=\frac{1}{2}$. Strongly convex functions satisfy the PL inequality [Braun2022-ew, Lemma 2.13]. Note that we will not have the primal gap bound under the HEB condition and weak convexity because an inequality like (B.5) does not hold from the HEB condition.
B.3 Scaling Inequalities, Geometric Hölder Error Bounds, and Contractions
@@
Scaling inequalities are a key tool in establishing convergence rates of FW algorithms. We will introduce two such inequalities that we will use in the following. The first scaling inequality is useful for analyzing the case where the optimal solution lies in the relative interior of the feasible region. While its formulation in [Braun2022-ew, Proposition 2.16] required $L$-smoothness of $f$, it is actually not used in the proof, and the results hold more broadly. We restate it here for the sake of completeness.
Proposition B.7 (Scaling inequality from convexity when $x^{*}\in\operatorname{int}P$ [Braun2022-ew, Proposition 2.16]).
@@
Proposition B.8 (Scaling inequality from weak convexity when $x^{*}\in\operatorname{int}P$).
@@
$\langle\nabla f(x),v\rangle\leq\langle\nabla f(x),x^{*}-rz\rangle=\langle\nabla f(x),x^{*}\rangle-r\|\nabla f(x)\|.$ (B.7)
@@
By rearranging (B.7) and using $\langle\nabla f(x),x-x^{*}\rangle\geq 0$, we obtain
@@
The authors in [Lacoste-Julien2015-gb] defined a geometric distance-like constant of a polytope, known as the pyramidal width, to analyze the convergence of the away-step Frank–Wolfe algorithm (and other variants that use the pyramidal width) over polytopes. It can be interpreted as the minimal $\delta>0$ satisfying the following scaling inequality (B.8), which plays a central role in establishing convergence rates for the away-step FW algorithm; see [Braun2022-ew] for an in-depth discussion.
Lemma B.9 (Scaling inequality via pyramidal width [Braun2022-ew, Theorem 2.26][Lacoste-Julien2015-gb, Theorem 3]).
@@
$\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle\geq\delta\frac{\langle\psi,x-y\rangle}{\|x-y\|}.$ (B.8)
Note that Lemma B.9 does not require the convexity of $f$, and we will use it for nonconvex optimization. When $\phi$ is $\sigma$-strongly convex and $\langle\psi,x-y\rangle\geq 0$, Lemma B.9 implies
@@
We will now introduce the more compact notion of geometric Hölder error bound condition, which simply combines the pyramidal width and the HEB condition of the function $f$; see also [Braun2022-ew, Lemma 2.27] for details when $q=2$.
Lemma B.10 (Geometric Hölder Error Bound).
@@
Combining (B.8) with Lemma B.3 for $\psi=\nabla f(x)$, we have
@@
Lemma B.11 (From contractions to convergence rates [Braun2022-ew, Lemma 2.21]).
@@
Appendix C Proof of Convergence Analysis for Convex Optimization
We would also like to stress that while we formulate some of the results for the case that $x^{*}\in\operatorname{int}P$, the results can be extended to the case that $x^{*}\in\operatorname{ri}P$, i.e., the relative interior of $P$, which we did not do for the sake of clarity. Basically, the analysis is performed in the affine space spanned by the optimal face of $P$ in that case; the interested reader is referred to [Braun2022-ew] for details on how to extend the results.
C.1 Sublinear Convergence
We establish a sublinear convergence rate of the FW algorithm under the smooth adaptable property. It is a similar result to [Vyguzov2025-ha, Theorem 1]. We conducted its proof following [Braun2022-ew, Jaggi2013-oz, Pokutta2024-qu]. The FW algorithm uses the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$ in the following theorem.
Theorem C.1 (Primal convergence of the Frank–Wolfe algorithm).

### v4

@@ B.2 Properties of Smooth Adaptable Property
Lemma B.4 (Extended Descent Lemma (Bolte et al., 2018, Lemma 2.1)).
@@
From this, it can be seen that the $L$-smad property for $(f,\phi)$ provides upper and lower approximations for $f$ majorized by $\phi$ with $L>0$. In addition, if $\phi(x)=\frac{1}{2}\|x\|^{2}$ on $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$, Lemma B.4 corresponds to the classical descent lemma. While the $L$-smad property might seem unfamiliar at first, it is a natural generalization of $L$-smoothness, and examples of functions $f$ and $\phi$ satisfying the $L$-smad property are given, e.g., in (Bauschke et al., 2017, Lemmas 7 and 8)(Bolte et al., 2018, Lemma 5.1), (Dragomir et al., 2021, Propositions 2.1 and 2.3), (Mukkamala and Ochs, 2019, Proposition 2.1), (Takahashi and Takeda, 2025, Proposition 24), (Takahashi et al., 2023, Theorem 1), and (Takahashi et al., 2026, Theorem 15).
Example B.5.
Other examples can be found in Sections 6.16.2, Appendices F.1F.4, and F.5.
  1. (i)
    Blind deconvolution: Let $f:\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\to\mathbb{R}$ be $f(h,x)=\frac{1}{4}\|Bh\odot Ax-y\|^{2}$, where $\odot$ denotes the Hadamard product (also known as the element-wise product). The function $f$ can be decomposed into a DC function $f(h,x)=f_{1}(h,x)-f_{2}(h,x)$ with $f_{1}(h,x)=\frac{1}{4}\|Bh\|_{4}^{4}+\frac{1}{4}\|Ax\|_{4}^{4}+\frac{1}{2}\left(\|Bh\odot Ax\|^{2}+\|y\odot Bh\|^{2}+\|Ax\|^{2}+\|y\|^{2}\right)$ and $f_{2}(h,x)=\frac{1}{4}\|Bh\|_{4}^{4}+\frac{1}{4}\|Ax\|_{4}^{4}+\frac{1}{2}\|y\odot Bh+Ax\|^{2}$, where $B\in\mathbb{R}^{m\times d_{1}}$, $A\in\mathbb{R}^{m\times d_{2}}$, and $y\in\mathbb{R}^{m}$. Let $\phi(h,x)=\frac{1}{4}(\|h\|^{2}+\|x\|^{2})^{2}+\frac{1}{2}(\|h\|^{2}+\|x\|^{2})$ For any $L$ satisfying
    $L\geq\sum_{j=1}^{m}\left(3\|b_{j}\|^{4}+3\|a_{j}\|^{4}+\|b_{j}\|\|a_{j}\|+|y_{j}|\|b_{j}\|^{2}+\|a_{j}\|^{2}\right),$
    where $b_{j}$ and $a_{j}$ are the $j$th row vectors of $B$ and $A$, respectively, $(f_{1},\phi)$ is $L$-smad (Takahashi et al., 2023, Theorem 1). Note that Bregman proximal DC algorithms require the smooth adaptable property of $(f_{1},\phi)$, instead of $(f,\phi)$.
  2. (ii)
    KL-NMF: Let $f:\mathbb{R}_{+}^{m\times r}\times\mathbb{R}_{+}^{r\times n}\to\mathbb{R}$ be $f(W,H)=\sum_{i=1}^{m}\sum_{j=1}^{n}(X_{ij}\log\frac{X_{ij}}{(WH)_{ij}}-X_{ij}+(WH)_{ij})$, where $X\in\mathbb{R}_{+}^{m\times n}$. $f$ is bounded by $\hat{f}(W,H)=\sum_{i,j}(X_{ij}\log X_{ij}-X_{ij}\sum_{l=1}^{r}\alpha_{ilj}\log\frac{W_{il}H_{lj}}{\alpha_{ilj}}-X_{ij}+(WH)_{ij})$ with $\alpha_{ilj}\in[0,1]$, $i=1,\ldots,m$, $l=1,\ldots,r$, $j=1,\ldots,n$. Let $\phi(W,H)=\sum_{i,l}(-\log W_{il}+\frac{1}{2}W_{il}^{2})+\sum_{l,j}(-\log H_{lj}+\frac{1}{2}H_{lj}^{2})$. Then, for any $L_{k}>0$ satisfying
    $L\geq\max\left\{\max_{i,l}\left\{\sum_{j=1}^{n}\alpha_{ilj}X_{ij}\right\},\max_{l,j}\left\{\sum_{i=1}^{m}\alpha_{ilj}X_{ij}\right\},m,n\right\},$
    the pair $(\hat{f},\phi)$ is $L$-smad on $\mathbb{R}_{++}^{m\times r}\times\mathbb{R}_{++}^{r\times n}$ (Takahashi et al., 2026, Theorem 4.1). Note that the majorization-minimization BPGs (Takahashi et al., 2026) require the smooth adaptable property of $(\hat{f},\phi)$, instead of $(f,\phi)$.
Remark B.6.
In general, it is difficult to choose the best $\phi$ such that $(f,\phi)$ is $L$-smad. On the other hand, it is known how to choose a suitable $\phi$ empirically. An easy way to find $\phi$ is to use elementary functions similar to $f$. See also the above examples, Section 6, and Appendix F.
@@
Lemma B.7 (Primal progress from the smooth adaptable property).
@@
Using Lemma B.4 and substituting $x_{+}$ for $x$ and $x$ for $y$, we have
@@
It holds that $D_{\phi}(x_{+},x)\leq\gamma^{1+\nu}D_{\phi}(v,x)$ for some $\nu>0$ by Lemma B.3. Therefore, this provides the desired inequality. ∎
@@
Lemma B.8 (Progress lemma from the Bregman short step-size).
@@
Substituting $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$ with $x_{t},v_{t}\in P$ for $x_{+}$ in Lemma B.7, we have
@@
B.3 Proof of Theorem 3.3
Proof of Theorem 3.3.
By following the same argument as (Pedregosa et al., 2020, Theorem 1), we have a bound on estimating $L$ as $(1-\log\eta/\log\tau)(t+1)+\max\{\log(\tau L/L_{-1}),0\}/\log\tau$. Let $n_{i}$ be the number of evaluations needed to estimate $\nu$. We have $\nu=\beta^{n_{i}-1}$, which provides $n_{i}=1+\log\nu/\log\beta$. Therefore, we obtain $n_{t}\leq\max\{(1-\log\eta/\log\tau)(t+1)+\max\{\log(\tau L/L_{-1}),0\}/\log\tau,(1+\log\nu/\log\beta)(t+1)\}$
Appendix C Properties from Sharpness
C.1 Results from Weak Convexity
A function $f$ is $\rho$-weakly convex and differentiable if and only if, for all $x,y\in\operatorname{dom}f$,
@@ C.1 Results from Weak Convexity
Any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex (Vial, 1983, Proposition 4.11) and Wu (2007). Since the proof of this fact is omitted therein, we provide the argument for the sake of completeness below.
Proposition C.1.
@@
Let $g=\lambda_{\min}(\nabla^{2}f(\cdot))$, where $\lambda_{\min}(\cdot)$ is the smallest eigenvalue. The function $g$ is continuous because $\lambda_{\min}(\cdot)$ and $\nabla^{2}f(\cdot)$ are continuous. Therefore, there exists the minimum value of $g$ on $P$ due to the continuity of $g$ and the compactness of $P$. For $\rho=|\min_{x\in P}g(x)|$, we have $\nabla^{2}f(x)+\rho I\succeq O$ for any $x\in P$, which implies $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex on $P$, i.e., $f$ is $\rho$-weakly convex on $P$. ∎
Minimizing a nonconvex $\mathcal{C}^{2}$-function over a compact set $P$ is equivalent to minimizing a weakly convex function over $P$. However, it is important to note that when $\rho$ becomes large, the assumptions of algorithms may not be satisfied (see also Theorems E.3 and 5.4).
@@
Lemma C.2 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex $f$).
$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (C.1)
@@
C.2 Primal Gap Bound
We show in the following lemma that the HEB condition immediately provides a bound on the primal optimality gap, which will be useful for establishing convergence rates of the proposed algorithms as well as the Łojasiewicz inequality (Bolte et al., 2007; Łojasiewicz, 1963; 1965), provided $f$ is convex.
Lemma C.3 (Primal gap bound from Hölder error bound).
@@ C.2 Primal Gap Bound
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}},$ (C.2)
@@
Remark C.4.
With the remark from above, we can immediately relate the HEB condition to the Łojasiewicz inequality (Bolte et al., 2007; Łojasiewicz, 1963; 1965); the Łojasiewicz inequality is used to establish convergence analysis (Bolte et al., 2007). To this end, using Equation C.2, we estimate $\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\leq\frac{\|\nabla f(x)\|\|x-x^{*}\|}{\|x-x^{*}\|}=\|\nabla f(x)\|$ to then obtain the weaker condition:
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\|\nabla f(x)\|^{\frac{q}{q-1}}=c^{\frac{1}{\theta}}\|\nabla f(x)\|^{\frac{1}{\theta}},$ (C.3)
where $c=\left(\frac{q}{\mu}\right)^{\frac{1}{q}}$ and $\theta=\frac{q-1}{q}$. Inequality Equation C.3 is called the cc-Łojasiewicz inequality with $\theta\in[0,1)$. If $\mathcal{X}^{*}\subseteq\operatorname{ri}P$, then the two conditions are equivalent. However, if the optimal solution(s) are on the boundary of $P$ as is not infrequently the case, then the two conditions are not equivalent as $\|\nabla f(x)\|$ might not vanish for $x\in\mathcal{X}^{*}$, whereas $\langle\nabla f(x),x-x^{*}\rangle$ does, i.e., the HEB condition is tighter than the one induced by the Łojasiewicz inequality.
@@
Lemma C.5 (Primal gap bound from the quadratic growth).
Let $f$ be a $\rho$-weakly convex function that satisfies the local $\mu$-quadratic growth condition such that $\rho<\mu$. Let $x^{*}$ be the unique minimizer of $f$ over $P$ and let $\zeta>0$. Then, the following holds: for all $x\in[f\leq f^{*}+\zeta]\cap P$,
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$ (C.4)
@@
$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\langle\nabla f(x),x-x^{*}\rangle=\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\|x-x^{*}\|.$ (C.5)
@@
Remark C.6.
In the same vein as the discussion in Remark C.4, we can immediately relate the local $\mu$-quadratic growth condition with $\zeta>0$ to the local Polyak–Łojasiewicz (PL) inequality. Using Equation C.4 in Lemma C.5, for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\|\nabla f(x)\|^{2},$ (C.6)
which is equivalent to the PL inequality (Polyak, 1963), also called the gradient dominance property (Combettes and Pokutta, 2020), with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$. The PL inequality is also equivalent to the Łojasiewicz inequality with $\theta=\frac{1}{2}$. Strongly convex functions satisfy the PL inequality (Braun et al., 2025, Lemma 2.13). Note that we will not have the primal gap bound under the HEB condition and weak convexity because an inequality like Equation C.5 does not follow from the HEB condition.
C.3 Scaling Inequalities, Geometric Hölder Error Bounds, and Contractions
@@
Scaling inequalities are a key tool in establishing convergence rates of FW algorithms. We will introduce two such inequalities that we will use in the following. The first scaling inequality is useful for analyzing the case where the optimal solution lies in the relative interior of the feasible region. While its formulation in (Braun et al., 2025, Proposition 2.16) required $L$-smoothness of $f$, it is actually not used in the proof, and the results hold more broadly. We restate it here for the sake of completeness.
Proposition C.7 (Scaling inequality from convexity when $x^{*}\in\operatorname{int}P$ (Braun et al., 2025, Proposition 2.16)).
@@
Proposition C.8 (Scaling inequality from weak convexity when $x^{*}\in\operatorname{int}P$).
@@
$\langle\nabla f(x),v\rangle\leq\langle\nabla f(x),x^{*}-rz\rangle=\langle\nabla f(x),x^{*}\rangle-r\|\nabla f(x)\|.$ (C.7)
@@
By rearranging Equation C.7 and using $\langle\nabla f(x),x-x^{*}\rangle\geq 0$, we obtain
@@
Lacoste-Julien and Jaggi (2015) defined a geometric distance-like constant of a polytope, known as the pyramidal width, to analyze the convergence of the away-step Frank–Wolfe algorithm (and other variants that use the pyramidal width) over polytopes. It can be interpreted as the minimal $\delta>0$ satisfying the following scaling inequality Equation C.8, which plays a central role in establishing convergence rates for the away-step FW algorithm; see Braun et al. (2025) for an in-depth discussion.
Lemma C.9 (Scaling inequality via pyramidal width (Braun et al., 2025, Theorem 2.26)(Lacoste-Julien and Jaggi, 2015, Theorem 3)).
@@
$\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle\geq\delta\frac{\langle\psi,x-y\rangle}{\|x-y\|}.$ (C.8)
Note that Lemma C.9 does not require the convexity of $f$, and we will use it for nonconvex optimization. When $\phi$ is $\sigma$-strongly convex and $\langle\psi,x-y\rangle\geq 0$, Lemma C.9 implies
@@
We will now introduce the more compact notion of geometric Hölder error bound condition, which simply combines the pyramidal width and the HEB condition of the function $f$; see also (Braun et al., 2025, Lemma 2.27) for details when $q=2$.
Lemma C.10 (Geometric Hölder error bound).
@@
Combining Equation C.8 with Lemma C.3 for $\psi=\nabla f(x)$, we have
@@
Lemma C.11 (From contractions to convergence rates (Braun et al., 2025, Lemma 2.21)).
@@
Appendix D Proof of Convergence Analysis for Convex Optimization
We would also like to stress that while we formulate some of the results for the case that $x^{*}\in\operatorname{int}P$, the results can be extended to the case that $x^{*}\in\operatorname{ri}P$, i.e., the relative interior of $P$, which we did not do for the sake of clarity. Basically, the analysis is performed in the affine space spanned by the optimal face of $P$ in that case; the interested reader is referred to Braun et al. (2025) for details on how to extend the results.
D.1 Sublinear Convergence
We recall a key property for analyzing convergence rates in convex optimization.
Lemma D.1 (Primal gap, dual gap, and Frank–Wolfe gap for convex $f$ (Pokutta, 2024, Lemma 4.1)).
$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle.$ (D.1)
We establish a sublinear convergence rate of the FW algorithm under the smooth adaptable property. It is a similar result to (Vyguzov and Stonyakin, 2025, Theorem 1). We conducted its proof following Braun et al. (2025); Jaggi (2013); Pokutta (2024). The FW algorithm uses the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$ in the following theorem.
Theorem D.2 (Primal convergence of the Frank–Wolfe algorithm).